### v1

# Think Outside the Policy: In-Context Steered Policy Optimization

Abstract
Existing Reinforcement Learning from Verifiable Rewards (RLVR) methods, such as Group Relative Policy Optimization (GRPO), have achieved remarkable progress in improving the reasoning capabilities of Large Reasoning Models (LRMs). However, they exhibit limited exploration due to reliance on on-policy rollouts where confined to the current policy’s distribution, resulting in narrow trajectory diversity. Recent approaches attempt to expand policy coverage by incorporating trajectories generated from stronger expert models, yet this reliance increases computational cost and such advaned models are often inaccessible. To address these issues, we propose In-Context Steered Policy Optimization (ICPO), a unified framework that leverages the inherent in-context learning capability of LRMs to provide expert guidance using existing datasets. ICPO introduces Mixed-Policy GRPO with Implicit Expert Forcing, which expands exploration beyond the current policy distribution without requiring advanced LRM trajectories. To further stabilize optimization, ICPO integrates Expert Region Reject Sampling to filter unreliable off-policy trajectories and Annealed Expert-Bonus Reward Shaping to balance early expert guidance with later autonomous improvement. Results demonstrate that ICPO consistently enhances reinforcement learning performance and training stability on mathematical reasoning benchmarks, revealing a scalable and effective RLVR paradigm for LRMs.
Think Outside the Policy: In-Context Steered Policy Optimization
Hsiu-Yuan Huang1,2††thanks: Equal contribution. Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3††thanks: Corresponding author., Saiyong Yang3, Yunfang Wu1,222footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University 3LLM Department, Tencent ✉ huang.hsiuyuan@stu.pku.edu.cn, wuyf@pku.edu.cn
1 Introduction
Large Reasoning Models (LRMs) excel at solving complex mathematical problems, and Reinforcement Learning from Verifiable Rewards (RLVR) provides a scalable way to refine their reasoning through symbolic correctness signals. Yet, limited exploration under standard Group Relative Policy Optimization (GRPO) Shao et al. (2024) often hinders robust reasoning improvement.
To address these limitations, recent work has explored combining Supervised Fine-Tuning (SFT) with Reinforcement Learning (RL) to strengthen the exploration of LRMs. One research direction interleaves SFT and RL updates Ma et al. (2025), enabling SFT to improve high-difficulty problem-solving while RL refines mid- to low-difficulty behaviors. However, repeated switching between paradigms introduces instability and inefficient convergence. Another line of work seeks to unify SFT and RL within a single training process through different strategies: incorporating SFT data as off-policy rollouts during RL Yan et al. (2025) to expand the exploration space; jointly optimizing SFT and RL objectives Fu et al. (2025) for tighter integration; and leveraging hints to bootstrap rollouts Liu et al. (2025); Zhang et al. (2025); Fu et al. (2025) to improve performance on harder prompts.
Together, these approaches reflect a growing consensus that combing SFT and RL can effectively expand the exploration of policy search into more promising reasoning spaces. However, as shown in Figure 1, three key challenges remain: (1) Exploration remains confined to the current policy distribution, as GRPO-based methods rely on on-policy sampling, resulting in limited trajectory diversity and might convergence to local optima, as detailed in Section 3. (2) Expanding the exploration space with trajectories from stronger LRMs incurs high computational cost and limited accessibility, since generating additional reasoning traces from advanced models is expensive and such models are not always available. (3) External trajectories are often noisy and unstable for training, incorporating them indiscriminately as off-policy rollouts can mislead policy updates and harm convergence stability.
To address these limitations, we propose In-Context Steered Policy Optimization (ICPO), a unified RL framework that exploits the LRM’s inherent In-Context Learning (ICL) capability to provide expert guidance instead of relying on external advanced LRMs. Specifically, (1) ICPO introduces Mixed-Policy GRPO with Implicit Expert Forcing (IEF), where expert-conditioned rollouts are generated through few-shot ICL guidance, enabling exploration beyond the current policy distribution and steering the model toward expert-aligned regions of the solution space. (2) To ensure reliable guidance, ICPO employs Expert Region Reject Sampling (ERRS), which filters out noisy or low-quality off-policy trajectories using verifiable reward signals, retaining only those that truly exhibit expert-aligned reasoning behavior and preventing misleading gradients from contaminating policy updates. (3) ICPO further incorporates an annealed expert bonus into the Reward Shaping (RS) design, enforcing strong expert-guided shaping in the early stage and gradually relaxing it to facilitate autonomous optimization as exploration capabilities grow.
Our experiments show that few-shot ICL rollouts exhibit superior quality and diversity, making them effective sources of expert guidance. Extensive experiments on multiple mathematical reasoning benchmarks further demonstrate that ICPO achieves substantial performance gains over vanilla and mixed-policy GRPO methods, with maximum average improvements of up to +4.17 and +2.79 points. On out-of-distribution benchmarks, ICPO also outperforms vanilla GRPO by up to +2.37 points.
Our contributions are three-fold:
We empirically show that few-shot ICL rollouts provide diverse and high-quality expert signals for mathematical reasoning task.
We propose In-Context Steered Policy Optimization, which include IEF that leverages the model’s inherent ICL ability without relying on stronger external LRMs, together with ERRS and RS for stable and efficient optimization.
ICPO delivers consistent improvements on mathematical reasoning benchmarks across model scales, demonstrating strong potential for LRM post-training.
2 Related Work
SFT and RL for LLM Reasoning.
GRPO-based RL improves reasoning by sampling and reinforcing high-quality trajectories, yet its exploration remains confined to the current policy distribution, leaving some prompts remain unsolved within feasible rollout budgets. Mainstream approaches introduce SFT to provide complementary trajectories, thereby expanding the reasoning space and enabling RL to further refine its behavior.
Several works explore hybrid SFT+RL strategies. ReLIFT Ma et al. (2025) alternates between RL and SFT by updating on failed rollouts. LUFFY Yan et al. (2025) incorporates SFT trajectories as off-policy samples using importance sampling. SRFT Fu et al. (2025) jointly optimizes SFT and RL objectives with an entropy-based weight on the SFT loss. Other approaches guide rollouts by concatenating partial SFT solutions as hints Liu et al. (2025); Zhang et al. (2025); Huang et al. (2025). However, these methods rely on advanced LRMs to supply SFT traces, which may incur additional computation overhead. In contrast, we leverage the inherent ICL ability of LRMs to obtain diverse rollouts from existing datasets, requiring neither external LRMs’ trajectories nor explicitly engineered hints.
3 Preliminary
Explicit Expert Forcing.
In traditional RL and RLHF, expert forcing explicitly constrains the policy to align with an expert policy , typically through imitation or KL-based regularization Hester et al. (2018); Haldar et al. (2023); Zhang et al. (2023); Hu et al. (2023). This explicit constraint stabilizes optimization and reduces reward variance, but it requires gradient-based imitation and access to an auxiliary expert model (e.g., a larger LRM), which can limit exploration in later stages.
Few-Shot ICL as Expert-Conditioned Inference.
Few-shot ICL provides a gradient-free way to inject expert priors through the input context. As illustrated in Figure 2 and Figure 3, we evaluate this effect from three complementary perspectives. (1) Accuracy: the 1-shot ICL setting consistently outperforms the 0-shot baseline, indicating that conditioning on demonstrations improves reasoning correctness. (2) Diversity: compared with temperature-based sampling, introducing 1-shot demonstrations expands the sampling space, yielding larger inter-trajectory similarity distances and enhanced exploratory diversity. (3) Distribution quality: under ICL-conditioned rollouts, the output distribution becomes more favorable—a higher proportion of previously incorrect generations are “flipped” to correct solutions compared with temperature perturbations, indicating that in-context steering provides a stronger and more targeted exploration signal. Taken together, these results support our view that few-shot ICL constitutes an effective expert-conditioned inference process.
From Few-Shot ICL to Implicit Expert Forcing.
Given expert demonstrations and a query , the model generates trajectories conditioned on:
Following the hypothesis-class view of ICL Hendel et al. (2023), the forward process of a Transformer can be decomposed into two functions:
where maps demonstrations to a task vector that encodes the expert behavior specific to that task Hendel et al. (2023); Li et al. (2023); Todd et al. (2024); Huang et al. (2024); Liu et al. (2024); Hojel et al. (2024), and represents the task-specific reasoning function that applies the to generate the prediction for query . This leads to the parametric representation:
indicating that ICL implicitly introduces an expert-induced prior that steers the rollout distribution toward expert-like regions—without any explicit optimization on . While few-shot ICL itself is an inference-time mechanism that does not update model parameters, we incorporate its induced trajectories into RL training to form IEF.
Group Relative Policy Optimization (GRPO).
GRPO is an efficient On-Policy optimization algorithm tailored for RL in LLMs, where the advantages for each token are computed in a group-relative manner without requiring an additional critic model to estimate token values. Given a set of rollouts sampled from the old policy , the normalized advantage is computed by:
Analogous to PPO Schulman et al. (2017), the GRPO objective is formulated as:
where is the importance ratio, and is the clipping function for variance reduction. To prevent the learned policy from drifting too far from the reference model, we retain the KL regularization term in GRPO, which is jointly optimized to ensure training stability and maintain controllable policy updates.
By leveraging ICL-conditioned rollouts within a mixed-policy GRPO framework, our approach enables expert-guided exploration to directly participate in policy optimization, effectively realizing an In-Context Steered Policy Optimization process.
4 Method
Figure 4 illustrates the overall ICPO training framework and the ICPO training process is detailed in Algorithm 1.
4.1 Mixed-Policy GRPO with Implicit Expert Forcing
To incorporate expert-conditioned exploration into group rollouts, we follow Yan et al. (2025) and extend GRPO into a Mixed-Policy setting, where each group consists of on-policy trajectories and trajectories generate under IEF , such that . We can recompute the group-normalized advantage (as in Eq. 4) over the mixed rollout set as:
where and .
The objective of mixed-policy GRPO with IEF can be written as follows:
where is the expert-conditioned importance weight. The mixed objective balances exploitation within the current policy support and exploration toward expert-aligned regions that are unreachable by standard on-policy rollouts.
Unlike prior work Yan et al. (2025), which adopts a model-based off-policy scheme by relying on an additional advanced LRM to provide expert trajectories, our mixed-policy GRPO with IEF operates as an input-conditioned off-policy method. Here, in-context demonstrations steer the same policy away from its natural output distribution, producing rollouts that differ from standard on-policy samples , thus effectively forming an off-policy distribution induced by IEF.
4.2 Expert Region Reject Sampling
Building upon the expert-conditioned off-policy branch above, we further restrict updates to those trajectories that demonstrably improve model performance. We define an Expert Region as the subset of states where expert conditioning yields superior guidance, steering the policy beyond its native distribution. A rollout generated under expert conditioning is accepted into this region if its reward exceeds a predefined threshold :
where is set to by default.
To prevent low-quality expert-conditioned traces from biasing training, we define a reject sampling operator that selectively retains trajectories within the Expert Region. Formally, performs reject sampling by restricting the expectation to trajectories that fall within the expert region:
where denotes the per-trajectory contribution to the objective. This filtering ensures that only high-reward expert-conditioned rollouts contribute to policy updates. The final objective of ICPO then becomes:
where normalizes over all valid tokens. The shaping function follows prior work Yan et al. (2025) and is defined as , where by default. This shaping biases learning toward expert-induced improvements while encouraging exploration.
4.3 Reward Shaping with Annealed Expert Bonus
[…]
This verifiable reward has been shown to reliably lead to successful scaling of RL training.
To encourage early imitation of expert-conditioned behavior while avoiding long-term over-reliance, we add a step-annealed bonus only to trajectories that have correct answer and within the expert region :
where denotes a linear decay scheduler over the training step , and denotes the bonus weight (set to in our experiments).
5 Experimental Setup
Dataset.
We follow Yan et al. (2025) and adopt the OpenR1-Math-220k dataset as our main training corpus. This dataset comprises mathematical problems sourced from NuminaMath-1.5 Li et al. (2024) and reasoning traces generated by the advanced LRM DeepSeek-R1 DeepSeek-AI (2025). Specifically, we use the filtered subset111https://huggingface.co/datasets/Elliott/Openr1-Math-46k, which excludes generations exceeding 8192 tokens as well as those identified as incorrect by Math-Verify222https://github.com/huggingface/Math-Verify. The resulting dataset contains approximately 45k verified reasoning prompts.
For the ablation study, we additionally train on a simpler dataset, Skywork-OR1-RL-Data333https://huggingface.co/datasets/Skywork/OR1-RL-Data He et al. (2025), to verify the generalization ability of our IEF under different reasoning conditions. This dataset is annotated with difficulty levels predicted by DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025), and we use the subset with difficulty level = 1, which corresponds to the easiest reasoning problems, as our simplified training corpus for controlled comparison.
Implementation Details.
We use Qwen3-1.7B and Qwen3-8B Yang et al. (2025) as the base model and employ GRPO Shao et al. (2024) as our RL algorithm. The KL regularization term is retained, while the entropy coefficient is set to 0.0 to stabilize training. We use a rollout batch size of 128 and an update batch size of 64. We generate 8 rollout trajectories per prompt with a maximum sequence length of 8192 tokens. Specifically, for the on-policy baseline, we use 8 on-policy rollouts. For our mixed-policy GRPO, we follow previous work Yan et al. (2025) and use 1 off-policy and 7 on-policy rollouts to ensure comparability. For implicit expert forcing, we randomly sample demonstrations from the MATH Hendrycks et al. (2021) training set, which contains 7.5k mathematical problems paired with high-quality solutions. Rollout generation uses a temperature of 1.0, and rewards are computed using Math-Verify. All models are trained for optimization steps, and we report results using the final checkpoint. Details of our evaluation setting are provided in the Appendix B.
Baseline Methods.
For RLVR-based methods, we compare against the vanilla GRPO baseline, which does not incorporate any external expert knowledge. We also include LUFFY Yan et al. (2025), which leverages trajectories generated by advanced LRMs as off-policy rollouts.
6 Results and Analysis
6.1 Main Results
| Expert Domain | In-Distribution Benchmarks | Overall | |||||||
| Model | Method | MATH-500 | AIME24 | AIME25 | AMC23 | Minerva | Olympiad | Avg. (Impr.) | |
| Qwen3-1.7B | GRPO | 83.60 | 28.44 | 22.50 | 66.72 | 40.81 | 48.15 | 48.37 | |
| ICPO | 86.80 | 31.25 | 26.25 | 70.39 | 44.12 | 56.44 | 52.54 | (+4.17) | |
| ICPO | 87.20 | 28.96 | 26.56 | 70.00 | 42.65 | 52.74 | 51.35 | (+2.98) | |
| Qwen3-8B | GRPO | 91.00 | 54.79 | 38.54 | 83.75 | 50.74 | 62.37 | 63.53 | |
| ICPO | 92.00 | 55.21 | 43.65 | 86.95 | 51.10 | 65.19 | 65.68 | (+2.15) | |
| ICPO | 92.00 | 56.15 | 40.94 | 92.00 | 51.47 | 64.30 | 65.04 | (+1.51) | |
| Method | ARC | GPQA | MMLU | Avg. (Impr.) |
| Qwen3-1.7B | ||||
| GRPO | 88.31 | 34.34 | 54.43 | 59.03 |
| ICPO | 88.14 | 27.78 | 55.45 | 57.12 (-1.91) |
| ICPO | 87.71 | 36.36 | 54.95 | 59.67 (+0.64) |
| Qwen3-8B | ||||
| GRPO | 95.82 | 51.01 | 71.98 | 72.94 |
| ICPO | 95.48 | 55.05 | 72.30 | 74.28 (+1.34) |
| ICPO | 95.56 | 55.05 | 75.31 | 75.31 (+2.37) |
The main experiments include two variants of our proposed ICPO framework: ICPO, which operates without reward shaping (RS), and ICPO, which incorporates RS to further enhance expert-domain alignment. To better understand their optimization behavior against GRPO, we further visualize the reward dynamics over training steps across different datasets, as shown in Figure 5, where both ICPO variants consistently achieve higher rewards throughout training.
Overall Improvements.
Table 1 reports the in-distribution reasoning performance, where MATH-500 serves as the expert domain. Across both model scales, ICPO consistently outperforms the vanilla GRPO baseline. For the smaller Qwen3-1.7B model, ICPO achieves an average improvement of +2.98 points over GRPO, while ICPO further stabilizes optimization with a +4.17 point overall gain. A similar trend is observed for the larger Qwen3-8B model, where ICPO and ICPO yield +2.15 and +1.51 average improvements, respectively. These consistent gains across scales demonstrate that our ICPO effectively steers the policy toward more expert-aligned regions in the parameter space.
Effect of Reward Shaping.
The variant ICPO introduces RS to explicitly amplify the advantage of trajectories falling within the expert domain . This design encourages the model to allocate higher probability mass to expert-like reasoning trajectories. Consequently, ICPO achieves stronger performance on the expert domain dataset (MATH-500), yielding +3.6 point improvements for Qwen3-1.7B. The improvement is less pronounced in non-expert domains, suggesting that RS primarily enhances expert-domain alignment rather than generic exploration. We note that the performance of ICPO could be further enhanced with higher-quality expert data, which would allow more precise shaping of expert-domain rewards and stronger alignment effects.
Out-of-Distribution Evaluation.
To assess whether ICPO preserves the model’s general reasoning capabilities, we further evaluate it on several OOD benchmarks, as shown in Table 2. The results indicate that ICPO maintains competitive or superior generalization performance compared to GRPO across both model scales, especially on larger model. For Qwen3 1.7B and 8B, ICPO achieves the average improvement of +0.64 and +2.37 over GRPO, suggesting that moderate RS enhances domain-specific reasoning while preserving broad OOD generalization. We also note that smaller models exhibit less stable OOD generalization than larger ones, likely due to their limited reasoning capacity.
Comparison of Expert Guidance Sources.
| Method | A. | E. | R. | MATH | AIME24/25 | AMC | Mnrv. | Avg. |
| Direct Comparison of Expert Guidance Sources | ||||||||
| LUFFY | ✓ | ✗ | ✗ | 91.00 | 53.12 / 36.98 | 85.31 | 52.21 | 63.72 |
| ICPO∗ | ✗ | ✓ | ✗ | 89.60 | 55.21 / 41.67 | 85.16 | 53.31 | 64.99 |
| In-Context Steered Policy Optimization | ||||||||
| ICPO | ✗ | ✓ | ✓ | 92.00 | 55.21 / 43.65 | 86.95 | 51.10 | 65.78 |
| ICPO | ✗ | ✓ | ✓ | 92.00 | 56.15 / 40.94 | 92.00 | 51.47 | 66.51 |
A. = Advanced LRM Trajectory, E. = Existing Dataset, R. = Expert Region Reject Sampling (ERRS). ICPO∗ = ICPO w/o ERRS.
We compare ICPO with LUFFY Yan et al. (2025), which incorporates trajectories generated by advanced LRMs into off-policy GRPO, as shown in Table 3. ICPO∗, which removes ERRS and is thus directly comparable to LUFFY in its source of expert guidance, already surpasses LUFFY by an average of +1.27 points. This demonstrates that IEF can steer the model toward a better policy distribution by leveraging existing datasets as contextual guidance, eliminating the need for costly external LRM computation. Moreover, incorporating ERRS and RS further improves performance by +2.06 and +2.79 points over LUFFY, confirming their complementary benefits.
6.2 Ablation Study
| Variant | MATH | AIME24/25 | AMC | Mnrv. | Avg. |
| Qwen3-1.7B | |||||
| ICPO | 87.20 | 28.96 / 26.56 | 70.00 | 42.65 | 51.07 |
| - RS | 86.80 | 31.25 / 26.25 | 70.39 | 44.12 | 51.76 |
| - ERRS | 85.60 | 32.19 / 25.94 | 66.80 | 42.28 | 50.56 |
| - IEF (GRPO) | 83.60 | 28.44 / 22.50 | 66.72 | 40.81 | 48.41 |
| Qwen3-8B | |||||
| ICPO | 92.00 | 56.15 / 40.94 | 92.00 | 51.47 | 66.51 |
| - RS | 92.00 | 55.21 / 43.65 | 86.95 | 51.10 | 65.78 |
| - ERRS | 89.60 | 55.21 / 41.67 | 85.16 | 53.31 | 64.99 |
| - IEF (GRPO) | 91.00 | 54.79 / 38.54 | 83.75 | 50.74 | 63.76 |
Mnrv. = Minerva, RS = Reward Shaping, ERRS = Expert Region Reject Sampling, IEF = Implicit Expert Forcing.
Effect of Each Component.
Table 4 reports ablation results by progressively removing each component of ICPO. On Qwen3-8B, we observe that all components contribute positively to the final performance. Implicit Expert Forcing (IEF) provides the largest gain (+1.23) over vanilla GRPO by injecting expert-conditioned guidance during rollout generation, and further enhances exploration by actively participating in the mixed-policy GRPO updates. Expert Region Reject Sampling (ERRS) improves accuracy by filtering out invalid expert-region trajectories. Reward Shaping (RS) further stabilizes optimization and benefits robustness across benchmarks. Across both model sizes, removing any single component consistently degrades accuracy, demonstrating that the three components are complementary and jointly essential for maximizing the effectiveness of ICPO.
Generalization of IEF Across Difficulty Levels.
| Variant | MATH | AIME24/25 | AMC | Mnrv. | Avg. |
| Skywork-OR1-RL-Data | |||||
| GRPO | 83.00 | 25.10 / 22.19 | 66.87 | 42.28 | 47.89 |
| ICPO | 86.00 | 26.77 / 24.06 | 69.61 | 42.65 | 49.82 |
| OpenR1-Math-220k | |||||
| GRPO | 83.60 | 28.44 / 22.50 | 66.72 | 40.81 | 48.41 |
| ICPO | 86.80 | 31.25 / 26.25 | 70.39 | 44.12 | 51.76 |
As shown in Table 5, results verify that IEF is consistently beneficial across training datasets of different difficulty levels. On the simpler Skywork dataset, ICPO improves GRPO by +1.93 average points, while on the more challenging OpenR1-Math corpus, the gain further increases to +3.35 average points. This demonstrates that IEF not only enhances learning on complex reasoning traces but also generalizes effectively to settings with weaker supervision.
6.3 Analysis on Source of Improvements
To better understand where the improvements of ICPO originate, we analyze the training dynamics of Qwen3-8B in Figure 6.
We observe that introducing IEF into mixed-policy GRPO drives the policy to deviate further from the reference distribution than vanilla GRPO, reflected by clear rises in entropy and KL divergence, accompanied by longer responses (Figure 6(a)). Importantly, this deviation is not uncontrolled drift: expert-conditioned rollouts encourage the model to explore expert-aligned yet previously unreachable regions of the solution space. As shown by the sharp peak in Figure 6(b), several prompts initially classified as All-Pass (i.e., groups whose GRPO rollouts are entirely correct) under GRPO begin producing more diverse reasoning strategies and transition to the Some-Pass group—precisely aligned with the onset of increased entropy and KL-loss. This indicates that IEF promotes deeper exploration beyond the solutions GRPO already masters. Meanwhile, ICPO consistently yields fewer Zero-Pass prompts than GRPO throughout training, indicating that expert guidance helps the model solve cases that GRPO fails on, thereby improving overall performance.
These observations together highlight that ICPO improves both the coverage and quality of successful rollouts without harming performance on easier problems. Nonetheless, the training dynamics demonstrate that ICPO enhances controllable exploration toward expert regions, which ultimately yields higher reasoning performance.
7 Conclusion
We present ICPO, a unified RLVR framework that enhances reasoning without relying on external expert models. Leveraging the inherent ICL capability of LRMs, ICPO introduces mixed-policy GRPO with IEF, which constructs expert-conditioned rollouts from existing datasets, improving data utilization and expanding exploration beyond the current policy distribution. To ensure stable optimization, ICPO further integrates ERRS to eliminate noisy off-policy trajectories and adopts RS to facilitate a smooth transition from expert-guided imitation to autonomous optimization. Experiments show that ICPO consistently improves RL performance, highlighting its promise as a scalable and general post-training paradigm for LRMs.
Ethics Statement
Use of AI Assistants
We certify that any use of AI tools, including ChatGPT, was strictly limited to linguistic refinement such as improving grammar, clarity, and style. All substantive ideas, analyses, and arguments presented in this work originate from the authors or from properly cited prior research.
Appendix A Prompt Construction
A.1 Prompt and Respones Format
Here we provide the detailed prompt format used for RL training, 0-shot and few-shot inference.
Appendix B Evaluation Settings
We evaluate on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 Li et al. (2024), Minerva Lewkowycz et al. (2022), Olympiad He et al. (2024), and MATH-500 Hendrycks et al. (2021). For AIME24, AIME25, and AMC23, which have relatively small test sets, we report Avg@32, while for the other benchmarks we report Pass@1. To assess generalization beyond in-domain reasoning, we further test on three out-of-distribution (OOD) benchmarks: ARC-C Clark et al. (2018), GPQA-Diamond Rein et al. (2023), and MMLU-Pro Wang et al. (2024), with multiple-choice options shuffled to prevent contamination. All evaluations are conducted using the LIMO framework Ye et al. (2025). During inference, we follow Yan et al. (2025) and set the generation temperature to 0.6 with a maximum response length of 8192 tokens. For few-shot ICL evaluation, we randomly sample demonstrations from the MATH Hendrycks et al. (2021) training set using 5 different random seeds, and report the average performance across them.